# Unveiling the Blind Spots: A Critical Examination of Fairness in Autonomous Driving Systems

### v2

@@ 4.2. RQ2: Fairness in Different Scenarios
Table 8. RQ2.3: MR and EOD of each pedestrian detector under non-rainy and rainy weather conditions. The statistically significant unfairness results are shaded.
Detectors Gender Age Skin Tone (LS: Light Skin, DS: Dark Skin)
Non-rainy weather Rainy weather Non-rainy weather Rainy weather Non-rainy weather Rainy weather
MR Male MR Female EOD MR Male MR Female EOD MR Adult MR Child EOD MR Adult MR Child EOD MR LS MR DS EOD MR LS MR DS EOD
YOLOX 9.58% 8.95% 0.63% 8.33% 5.17% 3.16% 10.45% 39.95% -29.50% 9.40% 30.00% -20.60% 4.19% 3.47% 0.72% 16.94% 13.51% 3.42%
RetinaNet 10.62% 11.35% -0.74% 10.09% 8.05% 2.04% 12.49% 40.65% -28.16% 13.61% 40.00% -26.39% 6.56% 3.81% 2.74% 21.77% 13.51% 8.26%
Faster RCNN 3.68% 3.29% 0.39% 4.39% 2.87% 1.51% 4.36% 23.09% -18.73% 4.86% 30.00% -25.14% 4.59% 2.60% 1.99% 15.32% 10.81% 4.51%
Cascade RCNN 3.68% 3.48% 0.20% 4.39% 2.87% 1.51% 4.53% 23.79% -19.25% 4.38% 25.00% -20.62% 3.93% 2.60% 1.33% 16.13% 8.11% 8.02%
ALFNet 23.52% 22.92% 0.60% 42.54% 40.80% 1.74% 25.00% 47.34% -22.34% 45.06% 75.00% -29.94% 34.90% 36.92% -2.02% 66.13% 64.86% 1.26%
CSP 25.29% 24.35% 0.94% 50.44% 44.25% 6.19% 26.86% 44.11% -17.25% 48.78% 65.00% -16.22% 53.05% 57.02% -3.97% 79.03% 89.19% -10.16%
MGAN 24.63% 25.26% -0.62% 50.88% 50.57% 0.30% 26.49% 41.11% -14.62% 53.00% 65.00% -12.00% 44.68% 47.83% -3.15% 72.58% 78.38% -5.80%
PRNet 32.73% 31.64% 1.09% 51.75% 50.57% 1.18% 33.19% 53.58% -20.39% 53.32% 75.00% -21.68% 51.99% 52.86% -0.87% 77.42% 81.08% -3.66%
Overall 16.72% 16.41% 0.31% 27.85% 25.65% 2.20% 17.92% 39.20% -21.28% 29.05% 50.63% -21.57% 25.49% 25.89% -0.40% 45.67% 44.93% 0.73%
In terms of the overall detection performance of all pedestrian detectors, the miss rate for each demographic group increases in the rainy weather condition, possibly due to droplets covering the camera and disrupting the detectors. Regarding the fairness, we find that the miss rate difference increases between the children and adults (with EOD from -21.28% to -21.57%) and between males and females (with EOD from 0.31% to 2.20%), indicating a higher miss rate for the children and males from non-rainy to rainy conditions. However, for these increases, we just observe a statistically significant difference from children and adults, the other two attributes do not exhibit statistically significant difference. Therefore, it can be concluded that rainy weather conditions have a minor impact on the fairness of pedestrian detectors.
Ans. to RQ2.3: Rainy weather conditions appear to have a slight effect on the equitable functioning of pedestrian detectors. There is an observed uptick in the discrepancy of miss rates when comparing demographics such as children versus adults, males versus females, and dark-skinned groups versus light-skinned groups. However, these increases are relatively minimal.
@@ 5.1. Fairness-performance Trade-off
Many studies (Ge et al., 2022; Wick et al., 2019; Chen et al., 2023; Hort et al., 2021) have revealed that fairness improvement usually comes at the cost of machine learning performance (e.g., accuracy). Therefore, developers need to grapple with the challenge of optimizing ML performance without compromising fairness, encapsulating this tension as the “fairness-performance trade-off.”
Nonetheless, we have noticed that pedestrian detectors exhibiting greater fairness (i.e., lower EOD values) can achieve superior performance (i.e., lower overall miss rates) under some environmental situations. For example, in the results of RQ2.1, we find that during daytime compared to nighttime, the eight pedestrian detectors achieved better detection results, concurrently displaying reduced EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2, depicted in Figure 3, reveal that heightened contrast levels may lead to improved detection performance and fairness.
Our findings provide a positive contrast to the previous fairness-performance trade-off theory, implicating that “You CAN have your cake and eat it” in the context of pedestrian detection. The immediately actionable insight is the existence of an optimal balance for enhancing both fairness and performance: by adjusting the brightness and contrast conditions of the captured images.
@@ 5.2. Implications
Implication for researchers. The significant unfairness exhibited by current pedestrian detectors offers valuable research opportunities for researchers. 1) Fairness improvement via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific image editing techniques to adjust contrast and brightness of input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Fairness improvement via multi-objective optimization. Given our discovery that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. 3) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 4) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage fellow researchers to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
@@
Implication for developers. Fairness is a critical non-functional requirement for Web-based applications, but our study demonstrates the existence of significant bias in state-of-the-art pedestrian detectors. It is imperative for developers to prioritize their efforts in this domain. Addressing these concerns goes beyond enhancing the quality of autonomous driving systems; it also serves to safeguard autnomous drving companies from ethical, reputational, financial, and legal repercussions that may arise in the event of violations of anti-discrimination laws.
@@
Implication for policy makers. Governments have a role in raising awareness about the potential biases in autonomous driving systems. 1) Regulatory measures. Autonomous driving systems play a crucial role in ensuring human safety. Our results reveal that fairness issues exist in modern pedestrian detection models. Therefore, it is essential for policy makers to enact regulations and standards that safeguard the rights of all individuals and address these concerns appropriately. 2) Enhanced protection for vulnerable pedestrians. Policies should also aim to address safety issues for vulnerable pedestrians such as children, who are, as revealed by this study, more likely to be overlooked by the current detection models.
@@ 6. Conclusion
This paper presents the first comprehensive study on fairness of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors with respect to gender, age, and skin tone. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. However, we observe balanced detection performance for male and female, light-skin and dark-skin individuals. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our investigation demonstrates that the detection performance for children and females is further reduced under low-brightness and low-contrast conditions, exacerbating the unfairness related to age and gender. Overall, this study sheds light on the fairness issues faced by existing pedestrian detectors, emphasizing the importance of addressing bias related to age and skin tone. The insights gained can pave the way for more fair and unbiased autonomous driving systems in the future.

### v3

@@ 4.2. RQ2: Fairness in Different Scenarios
Table 9. (RQ2.3) Miss rates and EOD of each pedestrian detector under non-rainy and rainy weather conditions. Statistically significant unfairness results are shaded. We find that rainy weather does not largely impact fairness regarding gender, age, and skin tone.
Gender
Detectors Non-rainy weather Rainy weather
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
YOLOX 9.58% 8.95% 0.63% 8.33% 5.17% 3.16%
RetinaNet 10.62% 11.35% -0.74% 10.09% 8.05% 2.04%
Faster RCNN 3.68% 3.29% 0.39% 4.39% 2.87% 1.51%
Cascade RCNN 3.68% 3.48% 0.20% 4.39% 2.87% 1.51%
ALFNet 23.52% 22.92% 0.60% 42.54% 40.80% 1.74%
CSP 25.29% 24.35% 0.94% 50.44% 44.25% 6.19%
MGAN 24.63% 25.26% -0.62% 50.88% 50.57% 0.30%
PRNet 32.73% 31.64% 1.09% 51.75% 50.57% 1.18%
Average 16.72% 16.41% 0.31% 27.85% 25.65% 2.20%
Age
Detectors Non-rainy weather Rainy weather
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
YOLOX 10.45% 39.95% -29.50% 9.40% 30.00% -20.60%
RetinaNet 12.49% 40.65% -28.16% 13.61% 40.00% -26.39%
Faster RCNN 4.36% 23.09% -18.73% 4.86% 30.00% -25.14%
Cascade RCNN 4.53% 23.79% -19.25% 4.38% 25.00% -20.62%
ALFNet 25.00% 47.34% -22.34% 45.06% 75.00% -29.94%
CSP 26.86% 44.11% -17.25% 48.78% 65.00% -16.22%
MGAN 26.49% 41.11% -14.62% 53.00% 65.00% -12.00%
PRNet 33.19% 53.58% -20.39% 53.32% 75.00% -21.68%
Average 17.92% 39.20% -21.28% 29.05% 50.63% -21.57%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors Non-rainy weather Rainy weather
MR LS MR DS EOD (Skin) MR LS MR DS EOD (Skin)
YOLOX 4.19% 3.47% 0.72% 16.94% 13.51% 3.42%
RetinaNet 6.56% 3.81% 2.74% 21.77% 13.51% 8.26%
Faster RCNN 4.59% 2.60% 1.99% 15.32% 10.81% 4.51%
Cascade RCNN 3.93% 2.60% 1.33% 16.13% 8.11% 8.02%
ALFNet 34.90% 36.92% -2.02% 66.13% 64.86% 1.26%
CSP 53.05% 57.02% -3.97% 79.03% 89.19% -10.16%
MGAN 44.68% 47.83% -3.15% 72.58% 78.38% -5.80%
PRNet 51.99% 52.86% -0.87% 77.42% 81.08% -3.66%
Average 25.49% 25.89% -0.40% 45.67% 44.93% 0.73%
From Table 9, we observe that rainy weather may potentially mitigate bias in pedestrian detectors. Specifically, under non-rainy conditions, all eight detectors exhibit significant bias against children, while under rainy conditions, three out of four pedestrian-specific detectors no longer show significant age bias any more. Similarly, under non-rainy conditions, two of eight detectors display significant bias against skin tone, whereas under rainy conditions, none do. However, the improvement in fairness due to rainy weather is marginal. The EOD difference between rainy and non-rainy conditions is 1.89% for gender, -0.29% for age, and 1.13% for skin tone.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.3: Rainy weather conditions decrease overall detection performance to a large extent but have a subtle impact on the fairness of pedestrian detectors. Specifically, three out of four pedestrian-specific detectors no longer exhibit significant age bias under rainy weather. However, the improvement in fairness is marginal.
@@ 5.1. Fairness-performance Trade-off
It is widely acknowledged that fairness improvement usually comes at the cost of machine learning performance (e.g., accuracy) (Ge et al., 2022; Wick et al., 2019; Chen et al., 2023; Hort et al., 2021). Therefore, developers need to grapple with the challenge of optimizing ML performance without compromising fairness, encapsulating this tension as the “fairness-performance trade-off.”
Nonetheless, we have observed that pedestrian detectors exhibiting greater fairness (i.e., lower absolute EOD values) can achieve superior performance (i.e., lower overall miss rates) under certain environmental conditions. For instance, in the results of RQ2.1, we find that during daytime compared to nighttime, the eight pedestrian detectors achieve better detection results while simultaneously displaying reduced EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2 reveal that the highest level of contrast consistently leads to improved fairness regarding gender, age, and skin tone, while also demonstrating the best overall detection performance in specific cases. These findings offer a positive counterpoint to previous fairness-performance trade-off trade-off theory, suggesting that we can enhance both detection performance and fairness by adjusting the brightness and contrast of captured images.
@@ 5.2.1. Implication for researchers
Our empirical results offer valuable research opportunities for researchers. 1) Fairness improvement of autonomous driving systems via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific post-processing image editing techniques to adjust contrast and brightness of captured input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Real-time adaptive fairness improvement of autonomous driving systems. Adapting software systems to increasingly dynamic environments poses significant challenges in software development, an area that has long been a important task in SE research (Qureshi and Perini, 2009). Our empirical findings contribute to this ongoing task by introducing a new challenge: designing real-time adaptive fairness improvement methods for autonomous driving systems. Specifically, our research reveals that autonomous driving systems exhibit greater biases under specific driving scenarios influenced by changes in brightness and contrast. Given the common occurrence of such changes in real-world driving conditions, there is an urgent need to develop specific methods to ensure the consistent fairness of autonomous driving systems in response to these environmental dynamics. 3) Fairness improvement of autonomous driving systems via multi-objective optimization. Given our finding that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. 4) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 5) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage fellow researchers to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
5.2.2. Implication for developers
@@
Fairness is a critical non-functional requirement for software applications, but our study demonstrates the existence of significant bias in state-of-the-art pedestrian detectors. It is imperative for developers to prioritize their efforts in this domain. Addressing these concerns goes beyond enhancing the quality of autonomous driving systems; it also serves to safeguard autonomous driving companies from ethical, reputational, financial, and legal repercussions that may arise in the event of violations of anti-discrimination laws. Moreover, it is crucial for developers of autonomous driving systems to consider the influence of various environmental factors on fairness during the development process.
5.2.3. Implication for policy makers
@@
Governments have a role in raising awareness about the potential biases in autonomous driving systems. 1) Regulatory measures. Autonomous driving systems play a crucial role in ensuring human safety. Our results reveal that fairness issues exist in modern pedestrian detection models. Therefore, it is essential for policy makers to enact regulations and standards that safeguard the rights of all individuals and address these concerns appropriately. 2) Enhanced protection for vulnerable pedestrians. Policies should also aim to address safety issues for vulnerable pedestrians such as children and females, who are, as revealed by this study, more likely to be overlooked by the current detection models.
5.3. Threats to Validity
Manual labeling. The labeling process involves possible subjectivity, posing threats to the validity of the analysis. To mitigate this threat, two annotators and an arbitrator are involved in the labeling process: first each image is independently labeled by two annotators, and in cases where conflicts in labeling arise, we seek the expertise of an arbitrator to resolve such discrepancies and arrive at a consensus. The inter-rater agreement between the two annotators is high, which demonstrates the reliability of the labeling schema and procedure adopted herein.
Selection of pedestrian detectors. Our study is based on eight pedestrian detectors, which may lead to possible selection bias in this study. To mitigate this threat, we select representative pedestrian detectors based on two considerations. On one hand, the selected detectors are widely-studied in pedestrian detection and autonomous driving literature (Cao et al., 2020; Zhang et al., 2016, 2017; Yu et al., 2020). On the other hand, the selected models cover the two typical types of pedestrian detection methods (i.e., one-stage detection and two-stage detection), ensuring a comprehensive representation of the techniques used in the field.
Selection of datasets. The choice of datasets presents a potential threat to validity. To address this concern, we use four real-world datasets extensively explored in autonomous driving research, including a total of 8,311 images. We augment the datasets with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. The extensive scale of data ensures the reliability of our results.
Selection of evaluation measures. The fairness measures that we employ may introduce potential limitations. To mitigate this concern, we consider three commonly-used fairness measures: SPD, EOD, and AOD. Upon careful discussion, we conclude that AOD may not be suitable for our study (Section 3.4), leading us to focus on SPD and EOD. We demonstrate that SPD and EOD lead to equivalent observations for pedestrian detection. Ultimately, we opt to use EOD (i.e., SPD) for our analysis, as it involves the comparison of the miss rates of different demographic groups. This ensures a consistent evaluation with the current pedestrian detection research, as miss rate is the most widely-adopted metric for measuring the performance of pedestrian detectors in the literature (Brandao, 2019; Kogure et al., 2022).
@@ 6. Conclusion
This paper presents the first comprehensive study on fairness of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors regarding gender, age, and skin tone. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. Additionally, pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Regarding skin tone, we observe balanced detection performance for light-skin and dark-skin individuals overall. As part of our contribution, we publicly release large-scale real-world pedestrian detection datasets with gender, age, and skin tone labels. These datasets aim to facilitate future fairness research in autonomous driving. The insights gained in this study can pave the way for more fair and unbiased autonomous driving systems in the future.
7. Data Availability
Our GitHub repository (git, 2024) contains datasets, sensitive attribute labels, scripts, and results of this work to facilitate replication and extension.